# Optimizing quantum violation for multipartite facet Bell inequalities

### v1

@@ Abstract
Nonlocality shapes quantum correlations, revealed through the violation of Bell inequalities. The intersection of all valid Bell inequalities is the so-called local polytope. In multipartite systems, characterizing the local polytope quickly becomes an intractable task as the system size increases. Optimizing Bell inequalities to maximize the ratio between their quantum value and classical bound is key to understanding multipartite nonlocality. We propose a gradient-based method for this optimization. Numerical results indicate that local maxima of this ratio typically correspond to facet Bell inequalities of the local polytope. This enables an iterative search for tight and robust Bell inequalities. Applied to permutation-invariant scenarios, the method provides tight Bell inequalities with large quantum violations and facilitates experimental certification of Bell correlations without full knowledge of the local polytope.
@@
Generalizing Bell nonlocality to the multipartite regime poses significant conceptual and practical challenges. As the number of parties increases, while the settings per party remain fixed, the number of correlation terms grows exponentially, and the complexity of characterizing the local polytope grows as $O((\exp N)^{(\exp N)})$ [21], rendering a complete characterization of multipartite nonlocality intractable  [22, 23, 24]. The local polytope, defined as the intersection of all valid Bell inequalities [25], can be characterized either by its vertices or its facets. Although generating the vertices is straightforward, the number grows exponentially with the number of parties. Deriving the corresponding facets quickly becomes computationally intractable [21, 26]. A practical approach is to restrict to few-body correlators and exploit symmetries, thereby greatly reducing the number of free coefficients in Bell inequalities and constraining the structure of correlations [27, 28, 29]. For instance, permutation symmetry among parties allows the construction of symmetric Bell inequalities that detect nonlocality in physically relevant many-body systems [30, 31], while remaining analytically and numerically tractable [27, 32]. Symmetry thus provides a promising route to construct scalable and experimentally relevant Bell inequalities for many-body systems [33, 34].
@@
In this Letter, we propose a gradient-based method to optimize Bell inequalities by enhancing the ratio between the quantum value for a fixed Hilbert space dimension, and the classical bound. We observe that the method converges to a local maximum of this ratio that corresponds to a tight Bell inequality. It simultaneously determines the classical bound, the measurement settings, and the quantum state achieving the quantum violation corresponding to such local maximum. We focus on the $(N,m,2)$ scenario, where each of the $N$ parties performs one of $m$ possible measurements with binary outcomes $\pm 1$. As a first benchmark, our optimization reproduces the CHSH inequality in the symmetrized $(2,2,2)$ scenario. We then apply it to the permutation-invariant (PI) Bell inequalities consisting of up to two-body correlators for general $(N,m,2)$ scenarios with $m=2$ and $3$. Our approach avoids enumerating all facets of the corresponding local polytope at once, a task that becomes intractable due to the exponential complexity with $N$. We identify families of tight Bell inequalities for $m=2$ and $3$, and obtain previously unknown tight Bell inequalities for $m=3$. Finally, we verify the maximum ratio on the affine projections and cross sections of both the primal and dual spaces, which reveals geometric insight into multipartite nonlocality.
@@
Setup.—A Bell inequality is given by a linear functional of the probability distribution $I=\sum_{\bm{a},\bm{x}}T_{\bm{a},\bm{x}}P(\bm{a}|\bm{x})\geq\beta_{C}$, where the coefficients $T_{\bm{a},\bm{x}}\in\mathbb{R}$ specify the inequality, $P(\bm{a}|\bm{x})$ denotes the conditional probability of obtaining outcomes $\bm{a}=(a_{1},\dots,a_{N})$ with $a_{i}=0$ or $1$ given measurement settings $\bm{x}=(x_{1},\dots,x_{N})$ with $x_{i}\in[m]=\{0,\dots,m-1\}$. The classical bound $\beta_{C}$ is defined as the minimum of $I$ over all local behaviors, $\beta_{C}=\min_{\vec{P}\in\mathcal{L}}I$, and $\vec{P}=\{P(\bm{a}|\bm{x})\}_{\bm{a},\bm{x}}$ [38], where $\mathcal{L}$ is the local polytope. The inequality is tight when the saturating local behaviors define a facet of the polytope $\mathcal{L}$, i.e., a supporting hyperplane of maximal dimension. We reformulate the linear functional in the correlator space as
$I=\sum_{\bm{\mu}}\alpha_{\bm{\mu}}\left\langle A_{\bm{\mu}}\right\rangle,$ (1)
@@
where the composite index $\bm{\mu}$ labels the $\ell$-body correlator $A_{\bm{\mu}}=A_{x_{i_{1}}}^{(i_{1})}\cdots A_{x_{i_{\ell}}}^{(i_{\ell})}$ acting on the sites $i_{1}$, $_{2},… ii_{\ell}$. The inputs on these sets are given by $\bm{x}_{\ell}=(x_{i_{1}},\dots,x_{i_{\ell}})\in[m]^{\ell}$. For binary outcomes, the correlator is defined by $\bigl\langle A_{\bm{\mu}}\bigr\rangle=\sum_{\bm{a}_{\ell}}(-1)^{\sum_{j=1}^{\ell}a_{i_{j}}}\,P(\bm{a}_{\ell}|\bm{x}_{\ell})$ with $\bm{a}_{\ell}=(a_{i_{1}},...,a_{i_{\ell}})$. The coefficient of each correlator is given by $\alpha_{\bm{\mu}}\in\mathbb{R}$ [28, 39, 40]. The classical bound of the Bell inequality can be rewritten in correlator form as $\beta_{C}=\min_{\left\langle A_{\bm{\mu}}\right\rangle\in\mathcal{L}}I$, where $\mathcal{L}$ now denotes the local polytope in the correlator space.
@@
Ratio optimization.—We define the ratio $\Delta:=\beta_{Q}/\beta_{C}$ that characterizes the potential of a Bell inequality to certify nonlocal correlations, with larger values indicating greater robustness to noise. By optimizing the coefficients $\alpha_{\bm{\mu}}$, we maximize the ratio $\Delta$, and our numerical results indicate that the local maxima typically correspond to facet Bell inequalities.
Note that $\beta_{Q}$ is a piecewise-smooth function with respect to the coefficients $\alpha_{\bm{\mu}}$, and its derivatives are given by [44]
@@
For the classical bound $\beta_{C}$, kinks occur when the set of saturating vertices changes on the local polytope. From the definition of $\beta_{C}$, at least one vertex saturates the Bell inequality. We fix the classical bound $\beta_{C}$ and iteratively add new saturating vertices determined by the gradient direction of $\beta_{Q}$. The coefficients $\alpha_{\bm{\mu}}$ are updated along this direction to increase the ratio
$\frac{\partial\Delta}{\partial\alpha_{\bm{\mu}}}=\frac{1}{\beta_{C}}\frac{\partial\beta_{Q}}{\partial\alpha_{\bm{\mu}}},$ (4)
subject to the constraints $\sum_{\bm{\mu}}\alpha_{\bm{\mu}}v_{\bm{\mu},j}=\beta_{C}$, where $v_{\bm{\mu},j}$ are the vertices saturating the classical bound. Other vertices yield larger classical values than the classical bound $\sum_{\bm{\mu}}\alpha_{\bm{\mu}}\left\langle A_{\bm{\mu}}\right\rangle\geq\beta_{C}$. By varying $\alpha_{\bm{\mu}}$, new vertices become active and are added to the constraints, progressively tightening the Bell inequality. A Bell inequality becomes tight when its hyperplane forms a facet of the local polytope, i.e., when it is saturated by a sufficient number of affinely independent vertices. We give the update rule of $\alpha_{\bm{\mu}}$ in [44].
The linear functional in Eq. (1) is defined with respect to the origin of correlator space, yet it can also be referenced to other points, e.g., $I=\sum_{\bm{\mu}}\alpha_{\bm{\mu}}(\left\langle A_{\bm{\mu}}\right\rangle-O_{\bm{\mu}})$ with some nonzero $O_{\bm{\mu}}$. Our method remains applicable as long as $O_{\bm{\mu}}$ lies within the local polytope, in which case the optimization identifies a nearby facet.
Figure 1: Illustration of the ratio-optimization procedure for a two-body Bell inequality. The maximal ratio identifies a tight Bell inequality. (a) Local (blue) polytope and quantum (red) convex set in the primal space, projected onto a two-dimensional plane spanned by the linear functionals $I_{1}$ and $I_{2}$. Blue dots represent the vertices of the local polytope after projection. (b) Corresponding objects in the dual-space cross-section, where cyan dots represent tight Bell inequalities. In both panels, the dashed curve indicates the tight inequality achieving the largest ratio. The light-green curve shows the projected trajectory of the coefficients $\alpha_{kl}$ during the optimization, plotted in the $(\lambda_{1},\lambda_{2})$ cross-section, demonstrating convergence to a CHSH inequality.
@@
Figure 1(a) shows a two-dimensional affine projection of the local polytope and the quantum convex set in the primal space of correlators, using the two linear functionals $I_{1}=(\langle A_{0}B_{0}\rangle+\langle A_{0}B_{1}\rangle+\langle A_{1}B_{0}\rangle-\langle A_{1}B_{1}\rangle)/2$ and $I_{2}=(-\langle A_{0}B_{0}\rangle+\langle A_{0}B_{1}\rangle+\langle A_{1}B_{0}\rangle+\langle A_{1}B_{1}\rangle)/2$. Figure 1(b) shows the corresponding dual-space cross-section (the space of Bell functionals up to scale). In general, dual extreme points are represented by coefficients $\lambda_{\bm{\mu}}=\alpha_{\bm{\mu}}/\beta$, so that the inequality reads $\sum_{\bm{\mu}}\lambda_{\bm{\mu}}\langle A_{\bm{\mu}}\rangle\geq-1$. Facet Bell inequalities correspond to vertices of the dual polytope at $\beta=\beta_{C}$; the dual quantum boundary is obtained at $\beta=\beta_{Q}$. We evaluate $\beta_{C}$ and $\beta_{Q}$ for the family $I=\lambda_{1}I_{1}+\lambda_{2}I_{2}$ spanned by $I_{1}$ and $I_{2}$. In the dual space, the admissible region is the set of $(\lambda_{1},\lambda_{2})$ such that every point in the primal space satisfies $\lambda_{1}I_{1}+\lambda_{2}I_{2}\geq-1$.
We then apply our gradient-based optimization to the coefficients $\alpha_{kl}$. The light-green curve shows the optimization trajectory (projected onto the dual cross-section), which converges to the CHSH facet, attaining the maximal quantum-to-classical ratio $\Delta^{\max}=\sqrt{2}$ ( $\beta_{Q}=-2\sqrt{2}$ vs. $\beta_{C}=-2$). The initial linear functional is $I=0.3\langle A_{0}B_{1}\rangle+0.3\langle A_{1}B_{0}\rangle-0.5\langle A_{1}B_{1}\rangle$, where the gradient vanishes; to escape this flat region we add small random perturbations to the coefficients $\alpha_{\bm{\mu}}$ [44].
@@
under the constraint $N+|\langle\mathcal{S}_{k}\rangle+\langle\mathcal{S}_{l}\rangle|\leq\langle\mathcal{Z}_{kl}\rangle\leq N-|\langle\mathcal{S}_{k}\rangle-\langle\mathcal{S}_{l}\rangle|$. To evaluate the classical bound $\beta_{C}$, we consider the local deterministic strategies for $\mathcal{S}_{k}$ and $\mathcal{Z}_{kl}$. It is known that vertices of the local polytope are characterized by $\mathcal{S}_{k}\in\{-N,-N+2,\ldots,N-2,N\}$ and $\mathcal{Z}_{kl}=-N+|\mathcal{S}_{k}+\mathcal{S}_{l}|$ or $N-|\mathcal{S}_{k}-\mathcal{S}_{l}|$ according to the sign of $\alpha_{kl}$ [27, 44]. For the quantum value, we adopt the symmetric measurement setting with the observables $\hat{\mathcal{S}}_{k}=2\cos\theta_{k}\hat{S}_{z}+2\sin\theta_{k}\hat{S}_{x}$ and $\hat{\mathcal{Z}}_{kl}=N\cos(\theta_{k}-\theta_{l})$, where $\hat{S}_{x}$ and $\hat{S}_{z}$ are the collective spin operators [27]. Restricting the measurements on qubits and onto the real circle of the Bloch sphere is sufficient to attain the maximum quantum violation for $m=2$ by virtue of Jordan’s lemma [45]. The quantum value is obtained as the variational minimum of the ground-state energy of the Hamiltonian $\hat{I}(\bm{\theta})$ over different measurement settings $\bm{\theta}=\{\theta_{k}\}_{k}$.
We apply our gradient-based optimization to the $(N,2,2)$ scenario [27]. For finite $N$, we seek the coefficients $\alpha_{\bm{\mu}}=(\alpha_{0},\alpha_{1},\alpha_{00},\alpha_{01},\alpha_{11})$ that maximize the ratio $\Delta_{N,2}$ and we denote the global maximum ratio for given $N$ as $\Delta_{N,2}^{\mathrm{max}}$. The vertices of the local polytope are the extreme points of the five-dimensional space spanned by $\langle\mathcal{S}_{\bm{\mu}}\rangle=(\langle\mathcal{S}_{0}\rangle,\langle\mathcal{S}_{1}\rangle,\langle\mathcal{S}_{00}\rangle,\langle\mathcal{S}_{01}\rangle,\langle\mathcal{S}_{11}\rangle)$. In this scenario, the numerical results show that the Bell inequalities become tight when the ratio attains a local maximum. Therefore, optimization of the coefficients can be used to find tight Bell inequalities. For example, for $N=3$, we obtain the tight Bell inequality with coefficients $\alpha_{\bm{\mu}}=(6,2,6,3,-2)$, which reaches the maximum ratio $\Delta_{3,2}^{\mathrm{max}}=1.11303$. For other values of $N$, the maximum ratios are summarized in Table LABEL:tab:finiteNwithm=2.
@@
$\alpha_{\bm{\mu}}=\alpha_{1,\bm{\mu}}+n\alpha_{2,\bm{\mu}},$ (6)
with $\alpha_{1,\bm{\mu}}=(1,-1,1,1,1)$ and $\alpha_{2,\bm{\mu}}=(1,1,0,0,0)$ and $n\in\{-2N+7,\,-2N+9,\,\ldots,\,2N-9,\,2N-7\}$. The maximum ratio is reached by $n=\pm 1$. Figure 2(a) and (b) show the affine projection and the cross section for this family of tight Bell inequalities for the $(12,2,2)$ scenario, respectively. The cyan dots represent the vertices of the local polytope. The blue and red dashed lines show the classical bound and the quantum value, respectively. This family includes the Bell inequality with coefficients $(2,0,1,1,1)$ which achieves the largest ratio as $N$ increases.
@@
NN $\Delta^{\mathrm{max}}_{N,2}$ $\beta_{C}$ $\alpha_{\bm{\mu}}$

### v2

@@ Abstract
Nonlocality shapes quantum correlations, revealed through the violation of Bell inequalities. The intersection of all valid Bell inequalities is the so-called local polytope. In multipartite systems, characterizing the local polytope quickly becomes an intractable task as the system size increases. Optimizing Bell inequalities to maximize the ratio between their quantum value and classical bound is key to understanding multipartite nonlocality. We propose a gradient-based method for this optimization. Numerical results indicate that local maxima of this ratio typically correspond to facet Bell inequalities of the local polytope. This enables an iterative search for tight and robust Bell inequalities. Applied to permutation-invariant scenarios, the method provides tight Bell inequalities with large quantum violations and facilitates experimental certification of Bell correlations without full knowledge of the local polytope. Moreover, analytical results of the maximum ratio are derived in the thermodynamic limit.
@@
Generalizing Bell nonlocality to the multipartite regime poses significant conceptual and practical challenges. As the number of parties increases, while the settings per party remain fixed, the number of correlation terms grows exponentially, and the complexity of characterizing the local polytope grows as $O((\exp N)^{(\exp N)})$ [21], rendering a complete characterization of multipartite nonlocality intractable  [22, 23, 24]. The local polytope, defined as the intersection of all valid Bell inequalities [25], can be characterized either by its vertices or its facets. Although generating the vertices is straightforward, the number grows exponentially with the number of parties. Deriving the corresponding facets quickly becomes computationally intractable [21, 26]. A practical approach is to restrict to few-body correlators and exploit symmetries, thereby greatly reducing the number of free coefficients in Bell inequalities and constraining the structure of correlations [27, 28]. For instance, permutation symmetry among parties allows the construction of symmetric Bell inequalities that detect nonlocality in physically relevant many-body systems [29, 30], while remaining analytically and numerically tractable [27, 31]. Symmetry thus provides a promising route to construct scalable and experimentally relevant Bell inequalities for many-body systems [32, 33].
@@
In this Letter, we propose a gradient-based method to optimize Bell inequalities by enhancing the ratio between the quantum value for a fixed Hilbert space dimension, and the classical bound. We observe that the method converges to a local maximum of this ratio that corresponds to a tight Bell inequality. It simultaneously determines the classical bound, the measurement settings, and the quantum state achieving the quantum violation corresponding to such local maximum. We focus on the $(N,m,2)$ scenario, where each of the $N$ parties performs one of $m$ possible measurements with binary outcomes $\pm 1$. As a first benchmark, our optimization reproduces the CHSH inequality in the symmetrized $(2,2,2)$ scenario. We then apply it to the permutation-invariant (PI) Bell inequalities consisting of up to two-body correlators for general $(N,m,2)$ scenarios with $m=2$ and $3$. Our approach avoids enumerating all facets of the corresponding local polytope at once, a task that becomes intractable due to the exponential complexity with $N$. We identify families of tight Bell inequalities for $m=2$ and $3$, and obtain previously unknown tight Bell inequalities for $m=3$. Finally, we verify the maximum ratio on the affine projections and cross sections of both the primal and dual spaces, which reveals geometric insight into multipartite nonlocality. We also obtain analytical results for the maximum ratio in the thermodynamic limit $N\rightarrow\infty$ for small $m$, while in the combined limit $N,m\to\infty$ it approaches $\coth(1)$.
@@
Setup.—A Bell inequality is given by a linear functional of the probability distribution $I=\sum_{\boldsymbol{a},\boldsymbol{x}}T_{\boldsymbol{a},\boldsymbol{x}}P(\boldsymbol{a}|\boldsymbol{x})\geq\beta_{C}$, where the coefficients $T_{\boldsymbol{a},\boldsymbol{x}}\in\mathbb{R}$ specify the inequality, $P(\boldsymbol{a}|\boldsymbol{x})$ denotes the conditional probability of obtaining outcomes $\boldsymbol{a}=(a_{1},\dots,a_{N})$ with $a_{i}=0$ or $1$ given measurement settings $\boldsymbol{x}=(x_{1},\dots,x_{N})$ with $x_{i}\in[m]=\{0,\dots,m-1\}$. The classical bound $\beta_{C}$ is defined as the minimum of $I$ over all local behaviors, $\beta_{C}=\min_{\vec{P}\in\mathcal{L}}I$, and $\vec{P}=\{P(\boldsymbol{a}|\boldsymbol{x})\}_{\boldsymbol{a},\boldsymbol{x}}$ [37], where $\mathcal{L}$ is the local polytope. The inequality is tight when the saturating local behaviors define a facet of the polytope $\mathcal{L}$, i.e., a supporting hyperplane of maximal dimension. We reformulate the linear functional in the correlator space as [20]
$I=\sum_{\boldsymbol{\mu}}\alpha_{\boldsymbol{\mu}}\left\langle A_{\boldsymbol{\mu}}\right\rangle,$ (1)
@@
where the composite index $\boldsymbol{\mu}$ labels the $\ell$-body correlator $A_{\boldsymbol{\mu}}=A_{x_{i_{1}}}^{(i_{1})}\cdots A_{x_{i_{\ell}}}^{(i_{\ell})}$ acting on the sites $i_{1},i_{2},\ldots,i_{\ell}$ with $1\leq i_{1}<i_{2}<\cdots<i_{\ell}\leq N$. The inputs on these sets are given by $\boldsymbol{x}_{\ell}=(x_{i_{1}},\dots,x_{i_{\ell}})\in[m]^{\ell}$. For binary outcomes, the correlator is defined by $\bigl\langle A_{\boldsymbol{\mu}}\bigr\rangle=\sum_{\boldsymbol{a}_{\ell}}(-1)^{\sum_{j=1}^{\ell}a_{i_{j}}}\,P(\boldsymbol{a}_{\ell}|\boldsymbol{x}_{\ell})$ with $\boldsymbol{a}_{\ell}=(a_{i_{1}},...,a_{i_{\ell}})$. The coefficient of each correlator is given by $\alpha_{\boldsymbol{\mu}}\in\mathbb{R}$. The classical bound of the Bell inequality can be rewritten in correlator form as $\beta_{C}=\min_{\left\langle A_{\boldsymbol{\mu}}\right\rangle\in\mathcal{L}}I$, where $\mathcal{L}$ now denotes the local polytope in the correlator space.
@@
Ratio optimization.—We define the ratio $\Delta:=\beta_{Q}/\beta_{C}$ that characterizes the potential of a Bell inequality to certify nonlocal correlations, with larger values indicating greater robustness to noise. By optimizing the coefficients $\alpha_{\boldsymbol{\mu}}$, we maximize the ratio $\Delta$, and our numerical results indicate that the local maxima typically correspond to facet Bell inequalities.
Note that $\beta_{Q}$ is a piecewise-smooth function with respect to the coefficients $\alpha_{\boldsymbol{\mu}}$, and its derivatives are given by [40]
@@
For the classical bound $\beta_{C}$, kinks occur when the set of saturating vertices changes on the local polytope. From the definition of $\beta_{C}$, at least one vertex saturates the Bell inequality. We fix the classical bound $\beta_{C}$ and iteratively add new saturating vertices determined by the gradient direction of $\beta_{Q}$. The coefficients $\alpha_{\boldsymbol{\mu}}$ are updated along this direction to increase the ratio
$\frac{\partial\Delta}{\partial\alpha_{\boldsymbol{\mu}}}=\frac{1}{\beta_{C}}\frac{\partial\beta_{Q}}{\partial\alpha_{\boldsymbol{\mu}}},$ (4)
subject to the constraints $\sum_{\boldsymbol{\mu}}\alpha_{\boldsymbol{\mu}}v_{\boldsymbol{\mu},j}=\beta_{C}$, where $v_{\boldsymbol{\mu},j}$ are the vertices saturating the classical bound. Other vertices yield larger classical values than the classical bound $\sum_{\boldsymbol{\mu}}\alpha_{\boldsymbol{\mu}}\left\langle A_{\boldsymbol{\mu}}\right\rangle\geq\beta_{C}$. By varying $\alpha_{\boldsymbol{\mu}}$, new vertices become active and are added to the constraints, progressively tightening the Bell inequality. A Bell inequality becomes tight when its hyperplane forms a facet of the local polytope, i.e., when it is saturated by a sufficient number of affinely independent vertices. We give the update rule of $\alpha_{\boldsymbol{\mu}}$ in [40].
The linear functional $I$ in Eq. (1) is defined with respect to the origin of correlator space, and we adopt the minimization convention. With this choice, any genuine quantum violation implies $\beta_{Q}/\beta_{C}>1$, reflecting that in the primal space the quantum set strictly contains the local polytope; our method also works in the maximization convention [20]. In addition, the functional can also be referenced to other points, e.g. $I=\sum_{\boldsymbol{\mu}}\alpha_{\boldsymbol{\mu}}\bigl(\langle A_{\boldsymbol{\mu}}\rangle-O_{\boldsymbol{\mu}}\bigr)$. Our method remains applicable as long as the point $O_{\boldsymbol{\mu}}$ lies inside the local polytope, in which case the optimization identifies a nearby facet.
Figure 1: Illustration of the ratio-optimization procedure for a two-body Bell inequality. The maximum ratio identifies a tight Bell inequality. (a) Local (blue) polytope and quantum (red) convex set in the primal space, projected onto a two-dimensional plane spanned by the linear functionals $I_{1}$ and $I_{2}$. Blue dots represent the vertices of the local polytope after projection. (b) Corresponding objects in the dual-space cross-section, where cyan dots represent tight Bell inequalities. In both panels, the dashed curve indicates the tight inequality achieving the largest ratio. The light-green curve shows the projected trajectory of the coefficients $\alpha_{kl}$ during the optimization, plotted in the $(\lambda_{1},\lambda_{2})$ cross-section, demonstrating convergence to a CHSH inequality.
@@
Figure 1(a) shows a two-dimensional affine projection of the local polytope and the quantum convex set in the primal space of correlators, using the two linear functionals $I_{1}=(\langle A_{0}B_{0}\rangle+\langle A_{0}B_{1}\rangle+\langle A_{1}B_{0}\rangle-\langle A_{1}B_{1}\rangle)/2$ and $I_{2}=(-\langle A_{0}B_{0}\rangle+\langle A_{0}B_{1}\rangle+\langle A_{1}B_{0}\rangle+\langle A_{1}B_{1}\rangle)/2$ [41, 37, 42]. Figure 1(b) shows the corresponding dual-space cross-section (the space of Bell functionals up to scale). In general, dual extreme points are represented by coefficients $\lambda_{\boldsymbol{\mu}}=\alpha_{\boldsymbol{\mu}}/\beta$, so that the inequality reads $\sum_{\boldsymbol{\mu}}\lambda_{\boldsymbol{\mu}}\langle A_{\boldsymbol{\mu}}\rangle\geq-1$. Facet Bell inequalities correspond to vertices of the dual polytope at $\beta=\beta_{C}$; the dual quantum boundary is obtained at $\beta=\beta_{Q}$. We evaluate $\beta_{C}$ and $\beta_{Q}$ for the family $I=\lambda_{1}I_{1}+\lambda_{2}I_{2}$ spanned by $I_{1}$ and $I_{2}$. In the dual space, the admissible region is the set of $(\lambda_{1},\lambda_{2})$ such that every point in the primal space satisfies $\lambda_{1}I_{1}+\lambda_{2}I_{2}\geq-1$.
We then apply our gradient-based optimization to the coefficients $\alpha_{kl}$. The light-green curve shows the optimization trajectory (projected onto the dual cross-section), which converges to the CHSH facet, attaining the maximum ratio $\Delta^{\max}=\sqrt{2}$ ( $\beta_{Q}=-2\sqrt{2}$ vs. $\beta_{C}=-2$). The initial linear functional is $I=0.3\langle A_{0}B_{1}\rangle+0.3\langle A_{1}B_{0}\rangle-0.5\langle A_{1}B_{1}\rangle$, where the gradient vanishes; to escape this flat region we add small random perturbations to the coefficients $\alpha_{\boldsymbol{\mu}}$ [40].
@@
under the constraint $N+|\langle\mathcal{S}_{k}\rangle+\langle\mathcal{S}_{l}\rangle|\leq\langle\mathcal{Z}_{kl}\rangle\leq N-|\langle\mathcal{S}_{k}\rangle-\langle\mathcal{S}_{l}\rangle|$. To evaluate the classical bound $\beta_{C}$, we consider the local deterministic strategies for $\mathcal{S}_{k}$ and $\mathcal{Z}_{kl}$. It is known that vertices of the local polytope are characterized by $\mathcal{S}_{k}\in\{-N,-N+2,\ldots,N-2,N\}$ and $\mathcal{Z}_{kl}=-N+|\mathcal{S}_{k}+\mathcal{S}_{l}|$ or $N-|\mathcal{S}_{k}-\mathcal{S}_{l}|$ according to the sign of $\alpha_{kl}$ [27, 40]. For the quantum value, we adopt the symmetric measurement setting with the observables $\hat{\mathcal{S}}_{k}=2\cos\theta_{k}\hat{S}_{z}+2\sin\theta_{k}\hat{S}_{x}$ and $\hat{\mathcal{Z}}_{kl}=N\cos(\theta_{k}-\theta_{l})$, where $\hat{S}_{x}$ and $\hat{S}_{z}$ are the collective spin operators [27]. Restricting the measurements on qubits and onto the real circle of the Bloch sphere is sufficient to attain the maximum quantum violation for $m=2$ by virtue of Jordan’s lemma [43]. The quantum value is obtained as the variational minimum of the ground-state energy of the Hamiltonian $\hat{I}(\boldsymbol{\theta})$ over different measurement settings $\boldsymbol{\theta}=\{\theta_{k}\}_{k}$.
We apply our gradient-based optimization to the $(N,2,2)$ scenario [27]. For finite $N$, we seek the coefficients $\alpha_{\boldsymbol{\mu}}=(\alpha_{0},\alpha_{1},\alpha_{00},\alpha_{01},\alpha_{11})$ that maximize the ratio $\Delta_{N,2}$ and we denote the global maximum ratio for given $N$ as $\Delta_{N,2}^{\mathrm{max}}$. The vertices of the local polytope are the extreme points of the five-dimensional space spanned by $\langle\mathcal{S}_{\boldsymbol{\mu}}\rangle=(\langle\mathcal{S}_{0}\rangle,\langle\mathcal{S}_{1}\rangle,\langle\mathcal{S}_{00}\rangle,\langle\mathcal{S}_{01}\rangle,\langle\mathcal{S}_{11}\rangle)$. In this scenario, the numerical results show that the Bell inequalities become tight when the ratio attains a local maximum. Therefore, optimization of the coefficients can be used to find tight Bell inequalities (see Table LABEL:tab:finiteNwithm=2).
@@
$\alpha_{\boldsymbol{\mu}}=\alpha_{1,\boldsymbol{\mu}}+n\alpha_{2,\boldsymbol{\mu}},$ (6)
with $\alpha_{1,\boldsymbol{\mu}}=(1,-1,1,1,1)$ and $\alpha_{2,\boldsymbol{\mu}}=(1,1,0,0,0)$ and $n\in\{-2N+7,\,-2N+9,\,\ldots,\,2N-9,\,2N-7\}$. The maximum ratio is reached by $n=\pm 1$. Figure 2(a) and (b) show the affine projection and the cross section for this family of tight Bell inequalities for the $(12,2,2)$ scenario, respectively. The cyan dots represent the vertices of the local polytope. The blue and red dashed lines show the classical bound and the quantum value, respectively. This family includes the Bell inequality with coefficients $(2,0,1,1,1)$ which achieves the largest ratio as $N$ increases.
@@
NN $\Delta^{\mathrm{max}}_{N,2}$ $\beta_{C}$ $\alpha_{\boldsymbol{\mu}}$